One number is seven less than a second number. Six times the first is 6 more than 4 times the second. Find the numbers
Answer by ikleyn(52802)

.

Let x be "the second number";

then the first number is  x-7, according to the condition.



Based on the second condition, you can write this equation

    6*(x-7) = 4*x + 6.



Simplify it and get the answer

    6x - 42 = 4x + 6

    6x - 4x = 6 + 42

      2x    = 48

       x    = 48/2 = 24.


ANSWER.  Second number is 24;  first number is seven less, i.i. 23-7 = 17.


CHECK.   Six times the first is 6*17 = 102;  4 times the second is 4*24 = 96.  The difference is 4.   ! Correct !

Solved.